The no. of solutions of dy/dx = (y+1) / (x -1)  , y(1) = 2 is


(a)none     (b)1      (c) 2        (d)infinite.

Grade:12

2 Answers

Badiuddin askIITians.ismu Expert
14 years ago

Dear Sanchit

dy/dx = (y+1) / (x -1)

dy /(y+1) =dx /(x-1)

ln(y+1) =ln(x-1) + ln c

y+1 =c(x-1)

Given function is not define at x=1

so no solution exits



Please feel free to post as many doubts on our discussion forum as you can.
 If you find any question Difficult to understand - post it here and we will get you the answer and detailed solution very quickly.
 We are all IITians and here to help you in your IIT JEE  & AIEEE preparation.

 All the best.
 
Regards,
Askiitians Experts
Badiuddin

Surya Anuraag Duvvuri
13 years ago

           dy/dx = (y+1)/(x-1)

          dy/(y+1) = dx/(x-1)

       Taking Integral on both sides,

       ln(y+1) = ln(x-1)+lnC

       Y(1)=2,

        Clearly at x=1 it does not exist

       therefore, no solutions exists.